how much of a 60% orange juice drink must be mixed with 4 gallons of a 30% orange juice drink to obtain a mixture of 50% orange juice?
Answer by mananth(16946)

percent ---------------- quantity
Orange juice I 30 ---------------- 4 gallons
Orange juice II 60 ---------------- x gallons
Mixture 50 ---------------- 4 + x gallons

30 * 4 + 60 x = 50 ( 4 + x )
120 + 60 x = 200 + 50 x
60 x + -50 x = 200 - -120
10 x = 80
/ 10
x = 8
8 gallons of 60 % Orange juice II II